REU 2005: REVISED SCHEDULE FOR WEEKS ONE AND TWO

ALL TALKS WILL BE IN ECKHART 206 (UNLESS THEY ARE IN RY251) FIRST WEEK: There will be an opening meeting Monday, June 20, at 9:00 in ECKHART 206. Pizza and coke will be available around noon that day, in the barn. Tutorials: 5:00 MWF, 5:30 TTH, evenings, and as arranged. Participate in the evenings – great time to meet friends and do math!!

Date Abert Babai Muchnik Fefferman Cattaneo
Mon, June 20 11:00–12:00  9:30–10:45 3:00–5:00

1:30–2:30

Tue, June 21 1:30–3:30 10:30–12:00 4:00–5:30 9:00–10:15
Wed, June 22 10:30–12:00 9:00–10:15 3:00–5:00

1:30–2:30

Thu, June 23  10:30–12:00 4:00–5:30 9:00–10:15

1:30–3:30

Fri, June 24 10:30–12:00  9:00–10:15 3:00–5:00

1:30–2:30

SECOND WEEK: YSP/SESAME preparation and training sessions take place Monday through Thursday, 9:00–12:00, June 27–July 1, for those teaching in YSP and SESAME. In addition, during the same time slot, Abert, Babai, May, and graduate students will teach a course on linear algebra and give other offerings for the apprentice program. Tutorial sessions will be arranged: lunch times, late afternoon, and evenings.

Date  YSP/SES Abert/Babai/May Abert  Babai  Bader Cattaneo
Apprentice  Problems Discrete
Mon, June 27 9:00–12:00  9:00–12:00  1:30–3:30  4:00–5:30
Tue, June 28 9:00–12:00  9:00–12:00  1:30–3:30  4:00–5:30
Wed, June 29 9:00–12:00  9:00–12:00  1:30–3:30 4:00–5:30
Thu, June 30 9:00–12:00  9:00–12:00  1:30–3:30 4:00–5:30

Fri, July 1

Here is the schedule in summary form: FIRST WEEK:

Abert (Problems) 11:00–12:00 and 1:30–2:30 M; 1:30–3:30 T; 10:30–12:00 and 1:30–2:30 F.

Babai (Discrete mathematics) 10:30–12:00 T; 10:30–12:00 and 1:30–2:30 W; 10:30–12:00 and 1:30–3:30 Th

Muchnik (Hyperbolic manifolds) 4:00–5:30 TTh Fefferman (Maximal functions) 9:30–10:45 M and 9:00–10:15 T–F Cattaneo (Transport phenomena) 3:00–5:00 MWF

SECOND WEEK:

YSP/SESAME Preparation 9:00–12:00 M-Th Abert/Babai/May (Apprentice) 9:00–12:00 M-Th Abert (Problems) 1:30–3:30 TTh Babai(Discrete mathematics) 1:30–3:30 M, 4:00–5:30 Th Bader (Hyperbolic manifolds) 4:00–5:30 T, 1:30–3:30 W Cattaneo (Transport phenomena) 4:00–5:30 MW